What iw the magnitude of the force that a 25x10^-6 C C charge exerts on a 10x10 ^ -6 C charge 8.5 cm away?


1
Expert's answer
2022-02-28T13:40:48-0500

The Coulomb's law says


"F=k\frac{q_1q_2}{r^2}\\
=9*10^9*\frac{25*10^{-6}*10*10^{-6}}{0.085^2}=311\;\rm N"